OpenStudy (cutiecomittee123):

Getting so tired of these but I need the practice and this is the last one so please help! 2x^2+y^2=33 x^2+y^2+2y=19

OpenStudy (perl):

Multiply the second equation by 2, then subtract the two equations. That will eliminate the x's

OpenStudy (cutiecomittee123):

Oh lol yeah solve the system of equations and then tell what the coordinates are of the solution that lies in quadrant 3

OpenStudy (cutiecomittee123):

wouldn't you multiply the second equation by -2?

OpenStudy (cutiecomittee123):

@perl

OpenStudy (freckles):

I thought we did this one together

OpenStudy (cutiecomittee123):

No they arent the same

OpenStudy (freckles):

they are exactly the same

OpenStudy (cutiecomittee123):

But I have another thing to ask and I dont feel like opening another question so can someone help? What conic section does this equation represent? 9x^2+4y^2+18x=16

OpenStudy (freckles):

how is this system 2x^2+y^2=33 x^2+y^2+2y=19 different from this system 2x^2+y^2=33 x^2+y^2+2y=19?

OpenStudy (cutiecomittee123):

Maybe they are the same but I really need the practice and the question was different

OpenStudy (freckles):

no it wasn't

OpenStudy (freckles):

but I'm done arguing

OpenStudy (perl):

You can multiply the second equation by -2